OpenStudy (anonymous):

Find sin A if sin4A=(2/3)

jimthompson5910 (jim_thompson5910):

Hints: sin(2x) = 2*sin(x)*cos(x) cos(2x) = 1 - 2*sin^2(x) and sin(4A) = sin(2A+2A) sin(4A) = sin(2A)cos(2A) + cos(2A)sin(2A) sin(4A) = 2*cos(2A)sin(2A)

jimthompson5910 (jim_thompson5910):

It's not complete, but I'll let you finish up

OpenStudy (anonymous):

i got that far. i'm just confused on the rest of the way.

OpenStudy (anonymous):

I got to 4sinAcosA(1-2sin^2A)=2/3. I just don't know how to convert this into a form for sin A.

jimthompson5910 (jim_thompson5910):

hmm let me think

jimthompson5910 (jim_thompson5910):

well it's really ugly, but I'll try my best to explain

jimthompson5910 (jim_thompson5910):

we know that sin(4A) = 2/3 so let's isolate A first

jimthompson5910 (jim_thompson5910):

sin(4A) = 2/3 4A = arcsin(2/3) A = (1/4)*arcsin(2/3) now we want to find sin(A), so apply the sine function to both sides to get sin(A) = sin((1/4)*arcsin(2/3))

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

yeah, it's been a while since i did arcsin but it should be easy for me to review. go on.

OpenStudy (anonymous):

wait arcsin is the inverse, isn't it? i'll try to solve the problem soon but if i have troubles i'll let you know. i'll put up my answer and steps regardless

jimthompson5910 (jim_thompson5910):

yes arcsine is the inverse of sine

jimthompson5910 (jim_thompson5910):

and from here, you will use the half angle identities which are cos(x/2) = sqrt( (1+cos(x))/2 ) sin(x/2) = sqrt( (1-cos(x))/2 )

OpenStudy (anonymous):

i appreciate it very much, thanks for the help

jimthompson5910 (jim_thompson5910):

yw